The system of equations may have a unique solution, an infinite number of solutions, or no solution.Use matrices to find the general solution of each system, if a solution exists. (If there is no solution, enter NO SOLUTION. If there are infinitely many solutions, express your answers in terms of z.

{{x+y+z=0}}}}
{{2x-y-z=0}}}
{{-x+2y+2z=0}}}
I have broken the matrix down to this:
[1 0 0 0]
[0 1 1 0]
[0 0 0 0]
How do i state this in terms of Z?

Answer by AnlytcPhil(1808)

The matrix

[1 0 0 | 0]
[0 1 1 | 0]
[0 0 0 | 0]

means

1x + 0y + 0z = 0
0x + 1y + 1z = 0
0x + 0y + 0z = 0

or

           x = 0
       y + z = 0

or

           x = 0
           y = -z

So the solution in terms of z is:

     (x,y,z) = (0,-z,z)

Edwin